OpenStudy (anonymous):

How do I solve this with elimination? 2x+3y=4 x+2y=-5

OpenStudy (anonymous):

Thank you :)

hero (hero):

2x+3y=4 x+2y=-5 y = (4-2x)/3 y = (-5-x)/2 Set y = y and cross multiply: 2(4-2x) = 3(-5-x) 8-4x = -15-3x x = 23 y = -14

hero (hero):

Sorry, I have to disagree with you on that

OpenStudy (anonymous):

both of our answers work checked with the calculator

hero (hero):

Weird, but okay

hero (hero):

I believe mine is the correct, correct solution though

OpenStudy (anonymous):

I'm so sorry, its not supposed to be two additions, the equation is; 2x+3y=4 x-2y=-5

OpenStudy (anonymous):

so its it supposed to be; x=-5-2y?

hero (hero):

c'mon man. You have to post the right equations. I did all that hard work for nothing

OpenStudy (anonymous):

i said sorry

hero (hero):

You're forgiven

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

moneybird, sorry but i didnt get where you got 2x-4y=-10 from?

hero (hero):

2x+3y=4 x-2y=-5 y = 4-2x)/3 y = (-5-x)/-2 Set y = y and cross multiply -2(4-2x) = 3(-5-x) -8+4x = -15 -3x 7x = -7 x = -1 y = 2

hero (hero):

Moneybird, whatever your approach is with these... It isn't working

hero (hero):

There's only one method that works everytime.

hero (hero):

By the way, moneybird...your solutions are not "true" solutions. They're the result of a flawed substitution method

OpenStudy (anonymous):

2x+3y=4 x-2y=-5 2x - 4y = -10 - 2x + 3y = 4 -7y = -14 y = 2 2x + 6 = 4 2x = -2 x = -1

OpenStudy (amistre64):

elimination gets us to the cramer rule \begin{array}l ax+by=n\\ cx+dy=m \end{array} \begin{array}r *-c)&ax+by=n\\ *a)&cx+dy=m \end{array} \begin{array}l -acx-cby&=-cn\\ \ \ \ acx+ady&=\ \ am\\ -----&----\\ \ \ \ \ y(ad-cb)&=am-cn \end{array} \[y=\frac{am-cn}{ad-cb}\] and x is similar to construct

hero (hero):

amistre...lol

OpenStudy (anonymous):

I saw my first mistake now sry

hero (hero):

Amistre always ends up doing some kind of extra cirriccular activity. Good catch money

OpenStudy (anonymous):

thank you guys for the help :)